OpenStudy (anonymous):

Adult IQ scores have a bell-shaped distribution with a mean of 100 and a standard deviation of 15. Use the Empirical Rule to find the percentage of adults with scores between 70 and 130.

OpenStudy (anonymous):

z(70)=70-100/15=-2 z(130)=130-100/15=2 area in the cure z(-2) to z(2) i.e 2 * area from P(0<Z<2)

OpenStudy (anonymous):

now u can find the region.

OpenStudy (anonymous):

alrite